OpenStudy (anonymous):

a1=2 and an=3(an-1)^2?? A 2,12,27 b 3,12,27 c 2,12,432 d 3,12,432

OpenStudy (anonymous):

@rvc can you help me with this??

OpenStudy (anonymous):

@Nnesha can you help???

OpenStudy (freckles):

so you have a_2 find a_2 by replace n with 2

OpenStudy (anonymous):

ok :) THANKYOU!!!

OpenStudy (freckles):

\[a_n=3(a_{n-1})^2 \\ a_2=3(a_{2-1})^2 \\ a_2=3(a_1)^2\\ a_2=3(2)^2 \\a_2=3(4) \\ a_2=12\] I will go ahead and give this one since all choices have a_2 is 12 then find a_3 by taking that same thing involving n and replacing n with 3

OpenStudy (xapproachesinfinity):

is that \(a_n=3a_{n-1}^2\)

OpenStudy (anonymous):

do you think its d or b

OpenStudy (xapproachesinfinity):

you didn't state what are you looking for by the way is a_2?

OpenStudy (anonymous):

the first 3 terms

OpenStudy (xapproachesinfinity):

all i see is a1 given to you and the the equation that generates the term of this sequence

OpenStudy (xapproachesinfinity):

from you choices it appears you are only looking for one term? are you sure that you are in need of 3 terms?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

then why the need of multiple choice answers?

OpenStudy (xapproachesinfinity):

I would ask you to recheck, did the question say find the 3th term of the sequence or find the first three terms?

OpenStudy (freckles):

\[a_1,a_2,a_3 \text{ are the first three terms } \] it sounds like they are asking you to state a_1 and a_2 anyways even though they are both given... you need to only really find the third term \[a_n=3 \cdot a_{n-1}^2\] have you replaced n with three yet ?

OpenStudy (xapproachesinfinity):

hmm where a_2 i don't see it

OpenStudy (freckles):

all the choices have a_2 is 12 technically it is given

OpenStudy (anonymous):

which represents the first 3 terms of the sequence? that the question with the prob

OpenStudy (xapproachesinfinity):

hmm it seems to me a though those given number are just one term? wouldn't you agree

OpenStudy (freckles):

a_1 is the first term a_2 is the second term a_3 is the third term there are three terms in each choice

OpenStudy (xapproachesinfinity):

hmm the choices say a) b) c) d) where do you see a1, a2,a3?

OpenStudy (freckles):

He said find the first three terms

OpenStudy (freckles):

Those terms are called a_1,a_2,a_3

OpenStudy (freckles):

a_1 and a_2 are already given

OpenStudy (freckles):

that is the first and second term

OpenStudy (freckles):

we just need to find the third term

OpenStudy (anonymous):

ohhh ok i got it its B you take the a2 into the equation and the 3 4 and you get B thankyou both for helping me :D:D

OpenStudy (xapproachesinfinity):

for the problem is miss worded and not stated correctly! but ate any rate finding the first three term just do as @freckles said

OpenStudy (freckles):

first of all @aswemayknow why are you saying the first term is 3 a_1 is given as 2 wasn't it? also can is you replace the n's with 3's and do the computation

OpenStudy (anonymous):

Thankyou :D:D

OpenStudy (xapproachesinfinity):

oh my bad! i realized my mistake heheh each choice is given 3 terms

OpenStudy (freckles):

\[a_n=3 \cdot a_{n-1}^2 \] put n as 3 \[a_3=3 \cdot a_{3-1} ^2 \] 3-1 is 2 what is a_2 ?

OpenStudy (xapproachesinfinity):

hehe B is not correct pal ssecond term is 12 first term is 2 given to you

OpenStudy (xapproachesinfinity):

B and D cannot be true since the first term is a1=2 for A and C since they both have a2=12 we need a3 to see which is true

OpenStudy (xapproachesinfinity):

use what @freckles did \(a_3=3a_2^2=2(12)^2\)

OpenStudy (freckles):

\[3(12)^2 \]

OpenStudy (xapproachesinfinity):

yep 3 I'm sloppy today no focus heheh